OpenStudy (anonymous):

A football player rushed 150 yards. Which measurement represents the same distance? 1500ft 15ft 450ft 50ft

OpenStudy (anonymous):

450ft

OpenStudy (anonymous):

1 yard is 3ft so 150*3=450 450ft

OpenStudy (anonymous):

1 yard is 3ft so,150 yard= 3*150=450 feet